OpenStudy (anonymous):

The function H described by H(x) = 2.75x + 71.48 can be used to predict the height in centimeters, of a woman whose humerus (the bone from the elbow to the sholder) is X cm long. Predict the height of a woman whose humreus is 40 cm long is?

OpenStudy (shubhamsrg):

what if i ask you value of 2x for x =4 ?

OpenStudy (ash2326):

H(x)=2.75x+71.48 We need to find the height of a woman whose humreus is 40 cm put x=40 \[H(40)=2.75\times 40+71.48\] Can you solve the rest?

OpenStudy (anonymous):

H=1.8558 @ash2326. I tried to find the value of h. Am I wrong?

OpenStudy (ash2326):

You're right but you wouldn't get 1.8558. Check your calculations

OpenStudy (anonymous):

If I round it up then it would be 2?

OpenStudy (ash2326):

\[2.75\times 40+71.48=110+71.48=181.48 cm\]

OpenStudy (anonymous):

Thank you, I just figured it out too with your help. Thank you.